OpenStudy (sashi786):

a person calculated the average of 10"three digit number ".but due to mistake he reversed the digits of a number and thus his average increased by 29.7 the difference between the unit digit and hundreds digit of that number is?? Plz anybody help me out with this qns

OpenStudy (amistre64):

hmm, im thinking of something like this to compare with \[\frac{K+a(100)+b(10)+c}{10}=M+29.7\] \[\frac{K+a+b(10)+c(100)}{10}=M\]

OpenStudy (amistre64):

since b(10) does not change, it can be absorbed into the K with no effect and you are left with a and c to play with

OpenStudy (amistre64):

\[100a + c = 10M + 297 - K\] \[100a + c-297 = 10M - K\] \[a + 100c = 10M - K\] .... \[100a+c-297=a + 100c\]

OpenStudy (amistre64):

\[99a-297=99c\] \[99a-99c=297\] \[99(a-c)=297\] looks to be something along those lines